OpenStudy (anonymous):

Rearrange this equation to isolate c a=b(1/c - 1/d)

OpenStudy (anonymous):

Solution: First, you need to apply the Distributive Property, a = b(1/c) + b(-1/d) a = b/c - b/d Next, we can move b/d to the other side: a + b/d = b/c - b/d + b/d a + b/d = b/c Next we need to cross multiply the expressions: (a+b/d)(c) = b Now, simply divide out the (a+b/d) from the left hand side: c = b/(a + b/d)

OpenStudy (anonymous):

I do not understand why you cross multiply and use c ?